INAF queue observing, 2017 Apr 18-19

Observer: R. Carini, M. Faccini
Telescope Operator: D. La Huerta
Support Astronomer: D. Thompson
Instrument: MODS; LUCI

Summary

At the sunset the sky is clear

Night duration (12\xB0 el twilight) 03:00am-11:50pm UT = 08 h 50m

Observing time: 8 hrs 50 m (100%)

Weather time loss: 0 hrs (0%)

technical time loss 0 hrs (0%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
Palazzi   ui, gr 3h completed
standard feige 34      
Grazian E ndwfsj05 dual grating 2 h completed
Grazian F sdss37 dual grating 1h 30m completed
Piconcelli P 1544+4120 HKspec 0.5 completed
Moscadelli I g074 H2 1h 15 m completed but some images with seeing more than 1.0''
Std FS149 H2    

Details

Dark for LUCI

dit=2.74, ndit=10, lir, int
  • luci1.20170419.0001-0005 *==luci1.20170419.0001-0005==
dit=6, ndit=10, lir, int
  • luci1.20170419.0006-0010 *==luci1.20170419.0006-0010==
dit=3, ndit=20, lir, int
  • luci1.20170419.0011-0015
  • luci1.20170419.0011-0015
dit=2.74, ndit=25, lir, int
  • luci1.20170419.0016-0020
  • luci1.20170419.0016-0020
dit=2.74, ndit=25, lir, int
  • luci1.20170419.0016-0020
  • luci1.20170419.0016-0020
dit=2.74, ndit=54, lir, int
  • luci1.20170419.0021-0025
  • luci1.20170419.0021-0025
dit=290, ndit=1, mer, nor
  • luci1.20170419.0026-0030
  • luci1.20170419.0026-0030
dit=180, ndit=1, mer, nor
  • luci1.20170419.0031-0035
  • luci1.20170419.0031-0035

skyflats with mods g u in blue channel, r i in red channel. In mods1 the flats are not good.
  • mods2r.20170419.0001-0004 good in r
  • mods2r.20170419.0009-0011 good in i
  • mods2b.20170419.0002-0003 good in u
  • mods2b.20179419.0007-0010 good in g

02:56 start Palazzi program with MODS execBinoMODS Atlas_img_1800.img
  • mods1r.20170419.0011-0015 ,mods1b.20170419.0011-0015
  • mods2r.20170419.0012-0016 ,mods2b.20170419.0011-0015
  • seeing about 1.0''
  • in these images there are the shadow of the probe, we change guide star.

03:26 start again the script, execBinoMODS Atlas_img_1800.img (g, r filetrs)
  • mods1r.20170419.0016-0017 ,mods1b.20170419.0016-0017 MODS1 is stuck after 2 images
  • mods2r.20170419.0017-0025 ,mods2b.20170419.0016-0024
  • 04:07 seeing about 0.9''

4:14 start new script Atlas_img_1800ui.img
  • mods1r.20170419.0018-0026 ,mods1b.20170419.0018-0027
  • mods2r.20170419.0026-0033 ,mods2b.20170419.0025-0033

05:00 start new script Atlas_img_1800.img
  • mods1r.20170419.0027-0035 ,mods1b.20170419.0028-0036
  • mods2r.20170419.0034-0042 ,mods2b.20170419.0034-0042

05:46 standard star Feige34
  • acqBinoMODS feige34.acq
  • mods1r.0036-0037
  • mods2r.0044-0046
  • execBinoMODS feige34.obs (dual grating)
  • mods1r.0038-0039, mods1b.0037-0038
  • mods2r.0047-0048, mods2b.0043-0044
06:16 To moves the telescope during the acquisition of the spectra for a misunderstanding. We have to take again the spectra.
  • mods1r.0040 , mods1b.0039, mods2r.0049, mods2b.0045... no good

06:19 acqMODS feige.acq
  • mods1r.0041-0042
  • mods2r.0050-0051

06:28 execBinoMODS feige34.obs
  • mods1r.0043-0044, mods1b.0040-0041
  • mods2r.0052-0053, mods2b.0046-0047

06:38 slew to Grazian program, target E
  • acqBinomods ndwf9j05_ls.acq
  • mods1r.0045-0048
  • mods2r.000054-0056

06:54 exeBinoMODS ndwf9j05_ls_4.obs
  • mods1r.0049-0061, mods1b.0042-0045
  • mods2r.0057-0060, mods2b.0048-0051
  • at the beginning of the acquisition of the spectra the seeing is about 0.7''
  • 7:19 seeing about 0.8''-0.9''
  • during the last image the seeing is not so good (aronu 1.2-1.4) we take another image

08:28 slew to Grazian program, target F
  • acqBinomods sdss37.acq
  • mods1r.0053-0055
  • mods2r.0062-0064 *seeing ~ 1

08:38 execBinoMODS* sdss37_ls_3.obs * mods1r.0056-0058, mods1b.0046-0048
  • mods2r.0065-0067, mods2b.0052-0054
  • mean seeing is 1.0'' with highest to 1.2'''
  • 09:07 seeing about 0.8''

09:20 switch to LUCI

09:53 Program Piconcelli, target P 1544+4120
  • luci1.20170419.0036-0040 , luci2.20170419.0036-0040 acquisition
  • luci1.20170419.0041-0044 , ==luci2.20170419.0041-0044==science
  • seeing about 0.8''-0.9''

10:31 telluric HIP71172
  • luci1.20170419.0045-0047 , luci2.20170419.0045-0047 acquisition
  • luci1.20170419.0048-0049 , ==luci2.20170419.0048-0049==science

10:45 Moscadelli I in H2
  • luci1.20170419.0050-0064 , luci2.20170419.0050-0064
  • from the second image seeing very variable between 0.9 to 1.6, avg seeing 1.3''
  • 11.06 avg seeing 1.2''
  • 11:08 avg seeing 1.1''
  • 11:10 avg seeing 1.06''
  • 11:30 seeing less than 1,0''
  • since bad seeing ad the beginning of the acquisition, we have taken two more images

11:36 we slew to the standard FS149 * luci1.20170419.0065-0068 , luci2.20170419.0065-0068

-- %USERSIG{RobertaCarini - 2017-04-19}%

Comments

 
Topic revision: r13 - 12 Jun 2017, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback